OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. 0 = x2 + 5x + 6 0 = x2 + 4x + 4 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 4x + 4 different from y = x2 + 5x + 6?

OpenStudy (anonymous):

very easy. firstly x^2=\[x ^{2}\] now your equation x^2+5x+6=0, we will solve by factorization althought we can also solve by quadratic formula i.e \[x=(-b \sqrt{b^2-4ac} )\div2a\] but we will solve by factorization because it is easy, x^2+5x+6=0 x^2+2x+3x+6=0 x(x+2)+3(x+2)=0 (x+2)(x+3)=0 either x+2=0 or x+3=0 either x=-2 or x=-3

OpenStudy (anonymous):

now 2nd equation x^2+4x+4 , we will factorize x^2+2x+2x+4=0 x(x+2)+2(x+2)=0 (x+2)(x+2)=0 x+2=0 or x+2=0 x=-2, the main difference between two equations is this that the first equation gave 2 answers and both satisfied the equation but 2nd equation gave two same answer and they obviously satisfied the equation

OpenStudy (anonymous):

these quadratic equation have something in common i.e they both can be solved by factorization and there is no need to apply quadratic formula